### v1

@@ 2.3 Computational costs
The training costs of our DL parametrization, namely CondensNet (i.e., the DL parametrization in our PCNN-GCM model) are relatively low: we spent a total of 40 GPU hours for training CondensNet, where we run 50 epochs for BasicNet (29 GPU hours), and 120 for ConCorrNet (11 GPU hours). BasicNet was constituted of 7 residual blocks (14 layers) and two separate ResMLP modules—one predicting $\mathrm{d}Q$ and one predicting $\mathrm{d}s$—each with 512 hidden neurons and ReLU activations, effectively yielding 1024 units in total. ConCorrNet was constituted of 6 residual blocks (12 layers) that included two 512-wide ResMLP modules employing sigmoid activations. Indeed, the overall DL parametrization is relatively small, with a total number of parameters equal to 1.7 million (1 million for BasicNet and 700,000 for ConCorrNet). For further details on the novel DL architecture and on training, the interested reader can refer to Methods section 4.1 and 4.1.2, respectively.
@@ 3 Discussion
This work introduces CondensNet, a novel DL parametrization, that embeds physical condensation constraints, to address long-term stability and accuracy issues. CondensNet is built on the important finding that the regulation of water vapor saturation is a significant factor affecting simulation stability in real-world land and ocean configurations. Unlike traditional super parametrization methods that can suppress anomalous water vapor transport, unconstrained DL parametrizations may fail to replicate moisture regulation, resulting in water vapor oversaturation. CondensNet directly learns saturation adjustment processes from super parametrization models (where in this work we used SPCAM), through a Condensation Correction Network (ConCorrNet), that works concurrently and adaptively with a residual multi-layer perceptron, namely BasicNet (see also Methods, section 4.1).
The resulting improvement in stability does not compromise predictive accuracy performance or the representation of energy and matter cycles (further results on this aspect are reported in Supplementary Information B). Moreover, the explicit inclusion of physical constraints enhances interpretability of the hybrid modeling, highlighting the crucial role played by the condensation process.
@@
In particular, we show how PCNN-GCM, through the novel CondensNet architecture, is able to run stable long-term climate simulations, following closely the SPCAM reference (Figure 2, in section 2.1), while maintaining physically-consistent results (Table 1 and Figure 3, in section 2.2; Supplementary Information section A). In addition, we show how the PCNN-GCM framework provides speedups in the order of 100x to 372x (depending on the number MPI ranks adopted), compared to the SPCAM reference (Figure 4, in section 2.3).
@@ 4 Methods
4.1 PCNN-GCM framework
4.1.1 CondensNet DL architecture and GCM model
CondensNet (Figure 5d) is a novel DL parametrization that learns and emulates the high-resolution cloud-resolving model (CRM) of SPCAM’s super parametrization khairoutdinov2001cloud (Figure 5b), where the atmospheric dynamics is driven by the Community Atmosphere Model version 5.2 (CAM5.2) neale2010description (Figure 5e), running at a horizontal resolution of $1.9^{\circ}\times 2.5^{\circ}$ with 30 vertical pressure levels, extending up to approximately 2 hPa. CAM5.2 is further coupled with the Community Land Model version 4.0 (CLM4.0) lawrence2011parametrization , using prescribed sea surface temperatures and sea ice concentrations according to the Atmospheric Model Intercomparison Project (AMIP) protocol.
@@ 4.1.1 CondensNet DL architecture and GCM model
Traditional GCMs like CAM use subgrid parametrization based on empirical models to represent cloud and convective processes (Figure 5a), which can introduce significant uncertainties. The GCM using super parametrization, like SPCAM khairoutdinov2001cloud , mitigates this issue by embedding a high-resolution cloud-resolving model (CRM) within each coarse grid cell (Figure 5b). In our study, the two-dimensional CRM of SPCAM consists of 32 grid points in the zonal direction and shares 30 vertical levels with the host model dynamics driven by CAM5.2. The host GCM includes all model components except for the parametrizations, namely: the dynamical core, the land model (CLM4.0), and the sea surface temperatures. Consequently, SPCAM, CAM5.2, and the hybrid modeling framework share identical host GCM components and simulation data coupling workflows.
The host GCM provides input variables including large-scale state variables such as water vapor $Q$, temperature $T$, surface pressure $P_{\mathrm{s}}$, and top-of-atmosphere solar insolation $Solin$. In addition, large-scale forcing variables such as water vapor forcing $\mathrm{d}Q_{\mathrm{l.s.}}$ and temperature forcing $\mathrm{d}T_{\mathrm{l.s.}}$ are supplied to further enhance the model’s predictive capability. CondensNet, our DL parametrization, inherits these input variables and returns predictions of water vapor tendency $\mathrm{d}Q$ and dry-static-energy tendency $\mathrm{d}s$ at each vertical level, along with predictions of direct and diffuse downwelling solar radiation fluxes to drive the coupled land surface model.
@@
  • Condensation Correction Network (ConCorrNet). ConCorrNet is designed to correct BasicNet’s predictions adaptively—thereby enforcing physical constraints associated with water vapor saturation—and, unlike methods that penalize non-physical predictions via the loss function or post-processing, it operates as an independent neural network that computes explicit correction terms and applies them selectively.
@@
In particular, we first employ a humidity detection module to identify grid points where the relative humidity ($rh$), defined with respect to liquid water, exceeds 100%. This process results in the creation of a humidity mask, ($\mathrm{Mask}_{\text{h}}$):
$\mathrm{Mask}_{\text{h}}(\mathrm{lon,lat,lev})$ =$=\begin{cases}1,&\text{if}\;rh>100\%\\[-5.0pt] 0,&\text{otherwise}\end{cases}$ (1)
where lon, lat, and lev represent the longitude, latitude, and vertical level indices of the grid points. We use $\mathrm{Mask}_{\text{h}}$ to mark regions where condensation is likely to occur; this marking directs ConCorrNet’s attention to these sensitive regions, but does not automatically enforce a correction. Whether a correction is applied depends on ConCorrNet’s adaptive learning from SPCAM data (i.e., the labels) and its subsequent predictions during simulation.
Next, ConCorrNet computes the condensation correction terms $\mathrm{d}Q_{\text{fix}}$ and $\mathrm{d}s_{\text{fix}}$ as follows:
$\mathrm{d}Q_{\text{fix}}$ $=\frac{Q_{v}-Q_{\mathrm{cond}}}{\Delta t}$ (2a)
$\mathrm{d}s_{\text{fix}}$ $=L_{v}\,\mathrm{d}Q_{\text{fix}}$ (2b)
where $Q_{v}$ is the predicted water vapor mixing ratio, $Q_{\mathrm{cond}}$ is the mixing ratio at condensation, $\Delta t$ is the simulation time step, and $L_{v}$ is the latent heat of condensation. These correction terms are then applied selectively using the humidity mask $\mathrm{Mask}_{\text{h}}$:
$\mathrm{d}Q_{\text{fixed}}$ $=\mathrm{d}Q-\mathrm{Mask}_{\text{h}}\odot\mathrm{d}Q_{\text{fix}}$ (3a)
$\mathrm{d}s_{\text{fixed}}$ $=\mathrm{d}s+\mathrm{Mask}_{\text{h}}\odot\mathrm{d}s_{\text{fix}},$ (3b)
where $\odot$ denotes element-wise multiplication. This active correction mechanism ensures that when the predicted water vapor mixing ratio exceeds the condensation threshold, the excess water vapor is appropriately condensed according to the saturation adjustment mechanism.
Water vapor condensation is triggered when the local relative humidity ($rh$) exceeds 100%, leading to the formation of liquid droplets or ice crystals. Generally $Q_{\mathrm{cond}}\geq Q^{\ast}$, where $Q^{\ast}$ is the saturation mixing ratio, that represents the water vapor content at this threshold and is commonly calculated as:
$Q^{\ast}=\frac{R_{d}\,e^{\ast}}{R_{v}\,(p-e^{\ast})}\approx\frac{0.622e^{\ast}}{p}.$ (4)
@@
In Equation 4, $R_{d}$ and $R_{v}$ are the specific gas constants for dry air and water vapor, respectively, $p$ is the atmospheric pressure, and $e^{\ast}$ is the saturation vapor pressure at temperature $T$, calculated using the Goff–Gratch equation goff1946low .
While ideally any excess water vapor above $Q^{\ast}$ would condense, in practice the process is influenced by additional factors such as aerosols and existing cloud particles.
Our CondCorrNet model of CondensNet adaptively ensures that when the water vapor mixing ratio exceeds the condensation threshold, the corresponding excess is condensed appropriately—with both the occurrence and the magnitude of condensation determined by the network’s learning of the labels, ensuring that the physical constraint is imposed in a data-driven, self-regulating manner.
4.1.2 Dataset and training details
@@
The output variables are the corresponding tendencies of moisture $\mathrm{d}Q$ and dry-static-energy $\mathrm{d}s$ at each vertical level (30 in total), as well as the four radiation fluxes ($SOLS$, $SOLL$, $SOLSD$, and $SOLLD$) in which reach to surface.
Notably, in CondensNet, following the traditional column-based parametrization design in GCMs, each neural network instance processes a single atmospheric column independently. During training, column samples from different spatial locations are randomly shuffled, as the network only needs to learn the vertical physical processes within individual columns. When coupled with the host GCM, while CondensNet instances operate independently for each grid point in each physics time step, columns exchange mass, momentum, and energy through dynamics, represented by the large-scale tendencies (input variables $\mathrm{d}Q_{\mathrm{l.s.}}$ and $\mathrm{d}T_{\mathrm{l.s.}}$). This design preserves both the original high parallel efficiency in parametrization and the essential horizontal coupling in dynamics.
Table 2 shows the input and output variables of CondensNet, including their physical dimensions and units. The inputs consist of vertical profiles of atmospheric state variables ($Q$, $T$), large-scale tendencies ($\mathrm{d}Q_{\mathrm{l.s.}}$, $\mathrm{d}T_{\mathrm{l.s.}}$), and surface conditions ($P_{\mathrm{s}}$, $Solin$). The outputs include physical tendencies of water vapor and dry-static-energy at each vertical level, along with surface radiation fluxes.
@@
Large-scale water vapor tendency $\mathrm{d}Q_{\mathrm{l.s.}}$ [kg/kg/s] 30 144 96
@@
Water vapor tendency $\mathrm{d}Q$ [kg/kg/s] 30 144 96
@@
The basic neural networks, BasicNet, is a pre-trained residual multilayer perceptron (i.e., that predicts basic tendencies of water vapor ($\mathrm{d}Q$) and dry-static-energy ($\mathrm{d}s$). It comprises 7 residual blocks (14 layers total) and 2 separate ResMLP modules—one for predicting $\mathrm{d}Q$ and one for $\mathrm{d}s$—each with 512 neurons in hidden layers and ReLU activations (effectively yielding 1024 neurons in total).
@@
The condensation correction network (ConCorrNet) is also a residual multilayer perceptron designed to adjust the predictions of BasicNet to enforce physical constraints related to water vapor saturation. ConCorrNet architecture includes 6 residual blocks, each containing 2 fully connected layers with 512 neurons, resulting in a total depth of 12 layers. We selected the sigmoid activation function based on its superior convergence performance observed in preliminary experiments.
During the fine-tuning of CondensNet, we freeze the parameters of the pre-trained BasicNet and train the ConCorrNet. Specifically, we use BasicNet configurations whose simulations crashed in order to test the condensation correction and ensure that the final predictions adhere to the physical constraints imposed. To this end, we defined two loss functions:
  1. 1.
    Overall Loss ($L_{\text{CondensNet}}$). This loss measures the difference between CondensNet final predictions and the true values from SPCAM simulation data
    $L_{\text{CondensNet}}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{\text{fixed}}-y_{\text{label}}\right)^{2}\end{aligned},$ (5)
    where $N$ is the number of samples, $y_{\text{label}}$ represents the true values, and $\hat{y}_{\text{fixed}}$ are the final predictions of CondensNet after applying the condensation correction, i.e., $\mathrm{d}Q_{\text{fixed}}$ and $\mathrm{d}s_{\text{fixed}}$ from Equation 3.
  2. 2.
    Condensation Correction Loss ($L_{\text{ConCorrNet}}$). This loss focuses on the condensation correction terms predicted by ConCorrNet:
@@
    $L_{\text{ConCorrNet}}=\frac{1}{N}\sum_{i=1}^{N}\text{Mask}_{\text{h}}\odot\left[\hat{y}_{\text{fix}}-(\hat{y}_{\text{fixed}}-y_{\text{label}})\right]^{2}\end{aligned},$ (6)
    where $\hat{y}_{\text{fix}}$ are the condensation correction terms, $\text{Mask}_{\text{h}}$ is a mask function applied to focus the loss on relevant regions, and $\odot$ denotes element-wise multiplication.
The Overall Loss updates the parameters of both ResMLP and ConCorrNet, ensuring that the combined model predictions match the SPCAM data. The Condensation Correction Loss updates only the parameters of ConCorrNet, allowing it to focus on learning the necessary oversaturation corrections without affecting the BasicNet. By optimizing both losses simultaneously, CondensNet effectively learns the necessary physical constraints, resulting in predictions consistent with its learning target (i.e., SPCAM data).
@@
4.2 Post-processing and error metrics
4.2.1 Means

### v2

@@ 2.3 Computational costs
The training costs of our DL parametrization, namely CondensNet (i.e., the DL parametrization in our PCNN-GCM model) are relatively low: we spent a total of 40 GPU hours for training CondensNet, where we run 50 epochs for BasicNet (29 GPU hours), and 120 for ConCorrNet (11 GPU hours). BasicNet is constituted of 7 residual blocks (14 hidden layers) and two separate Residual Multilayer Perceptron (ResMLP) neural networks – one predicting $\mathrm{d}Q$ and one predicting $\mathrm{d}s$. Each ResMLP has 512 neurons per hidden layer and uses ReLU activation functions. ConCorrNet is constituted of 6 residual blocks (12 hidden layers) and two separate ResMLP, one – one predicting $\mathrm{d}Q_{\mathrm{fix}}$ and one predicting $\mathrm{d}s_{\mathrm{fix}}$. Each ResMLP has 512 neurons per hidden layer and uses sigmoid activations. Indeed, the overall DL parametrization is relatively small, with a total number of parameters equal to 1.7 million (1 million for BasicNet and 700,000 for ConCorrNet). For further details on the novel DL architecture and on training, the interested reader can refer to Methods section 4 and 4.2, respectively.
@@
Performance metrics were obtained from CESM’s built-in timing diagnostics in CESM 1.1.1 (the ccsm_timing.\$CASE.\$date file under \$CASEROOT/timing/), which report simulated years per wall-day (SYPD), PE-hours per simulated year, seconds per model day, and per-component timings; we follow these standard CESM metrics throughout.
@@
For implementation details of the runtime coupling between the Fortran-based host and the Python-based DL parametrization, and for a discussion of scalability limitations and potential remedies, please refer to Supplementary Information section D.
@@ 3 Discussion
This work introduces CondensNet, a novel DL parametrization, that embeds physical condensation constraints, to address long-term stability and accuracy issues. CondensNet is built on the important finding that the regulation of water vapor saturation is a significant factor affecting simulation stability in real-world land and ocean configurations. Unlike traditional super parametrization methods that can suppress anomalous water vapor transport, unconstrained DL parametrizations may fail to replicate moisture regulation, resulting in water vapor oversaturation. CondensNet directly learns saturation adjustment processes from super parametrization models (where in this work we used SPCAM), through a Condensation Correction Network (ConCorrNet), that works concurrently and adaptively with a residual multi-layer perceptron, namely BasicNet (see also Methods, section 4).
The resulting improvement in stability does not compromise predictive accuracy performance or the representation of energy and matter cycles (further results on this aspect are reported in Supplementary Information C.1). Moreover, the explicit inclusion of physical constraints enhances interpretability of the hybrid modeling, highlighting the crucial role played by the condensation process.
@@
In particular, we show how PCNN-GCM, through the novel CondensNet architecture, is able to run stable long-term climate simulations, following closely the SPCAM reference (Figure 2, in section 2.1), while maintaining physically-consistent results (Table 1 and Figure 3, in section 2.2 as well as Supplementary Information section A). In addition, we show how the PCNN-GCM framework provides speedups in the order of 100x to 372x (depending on the number MPI ranks adopted), compared to the SPCAM reference (Figure 4, in section 2.3). PCNN-GCM is also faster and more accurate than CAM5 in certain aspects, such as precipitation and intraseasonal variability, and unlike CAM, supports substantial GPU acceleration. For example, when run on modest computational resources (a single 24-core node), PCNN-GCM is nearly four times faster than CAM5 with GPU acceleration.
@@ 4 Methods
4.1 CondensNet DL architecture and GCM model
CondensNet (Figure 5d) is a novel DL parametrization that learns and emulates the high-resolution cloud-resolving model (CRM) of SPCAM’s super parametrization khairoutdinov2001cloud (Figure 5b), where the atmospheric dynamics is driven by the Community Atmosphere Model version 5.2 (CAM5.2) neale2010description (Figure 5e), running at a horizontal resolution of $1.9^{\circ}\times 2.5^{\circ}$ with 30 vertical pressure levels, extending up to approximately 2 hPa, and employing a simulation timestep of 30 minutes. CAM5.2 is further coupled with the Community Land Model version 4.0 (CLM4.0) lawrence2011parametrization, using prescribed sea surface temperatures and sea ice concentrations according to the Atmospheric Model Intercomparison Project (AMIP) protocol.
@@ 4.1 CondensNet DL architecture and GCM model
Traditional GCMs like CAM use subgrid parametrization based on empirical models to represent cloud and convective processes (Figure 5a), which can introduce significant uncertainties. GCMs that use super parametrizations, like SPCAM khairoutdinov2001cloud, mitigate this issue by embedding a high-resolution cloud-resolving model (CRM) within each coarse grid cell (Figure 5b). In our study, the two-dimensional CRM of SPCAM consists of 32 grid points in the zonal direction and shares 30 vertical levels with the host model dynamics driven by CAM5.2. The host GCM includes all model components except for the parametrizations, namely: the dynamical core, the land model (CLM4.0), and the sea surface temperatures. Consequently, SPCAM, CAM5.2, and the hybrid modeling framework share identical host GCM components and simulation data coupling workflows.
The host GCM provides input variables including large-scale state variables such as water vapor $Q$, temperature $T$, surface pressure $P_{\mathrm{s}}$, and top-of-atmosphere solar insolation $Solin$. In addition, large-scale forcing variables such as water vapor forcing $\mathrm{d}Q_{\mathrm{l.s.}}$ and temperature forcing $\mathrm{d}T_{\mathrm{l.s.}}$ are supplied to further enhance the model’s predictive capability. CondensNet, our DL parametrization, inherits these input variables and returns predictions of water vapor tendency $\mathrm{d}Q$ and dry-static-energy tendency $\mathrm{d}s$ at each vertical level, using an independent ResMLP model from  wang2022stable) to predict downwelling solar radiation fluxes to drive the coupled land surface model.
@@
  • Condensation Correction Network (ConCorrNet). ConCorrNet is designed to predict physically-constrained corrections to BasicNet’s outputs. It is also an NN, operating as an independent, predictive module that is only activated adaptively in regions of unphysical atmospheric conditions (mainly oversaturation) to prevent model instabilities.
@@
In particular, we identify oversaturated grid points by comparing the prognostic specific humidity ($Q$) with the saturation specific humidity ($Q^{\ast}$). A grid point is marked for subsequent adaptive physical constraint if it satisfies the condition $Q>Q^{\ast}$, which is equivalent to a relative humidity exceeding 100%. This process results in the creation of a humidity mask, ($\mathrm{Mask}_{\text{h}}$):
$\mathrm{Mask}_{\text{h}}(\mathrm{lon,lat,lev})$ $=\begin{cases}1,&\text{if}\;Q>Q^{\ast}(rh>100\%)\\[-1.00006pt] 0,&\text{otherwise}\end{cases}$ (1)
where lon, lat, and lev represent the longitude, latitude, and vertical level indices of the grid points. The saturation specific humidity ($Q^{\ast}$) represents the maximum mass of water vapor that a unit mass of moist air can hold at a given temperature ($T$) and pressure ($p$). It is precisely defined and commonly approximated as:
$Q^{\ast}=\frac{\epsilon\cdot e^{\ast}}{p-(1-\epsilon)e^{\ast}}\approx\frac{0.622e^{\ast}}{p}$ (2)
@@
where, $p$ is the local atmospheric pressure, $e^{\ast}$ is the saturation vapor pressure at temperature $T$ (calculated using a formulation such as the Goff–Gratch equation goff1946low), and $\epsilon=R_{d}/R_{v}\approx 0.622$ is the ratio of the specific gas constants for dry air ($R_{d}$) and water vapor ($R_{v}$).
We use $\mathrm{Mask}_{\text{h}}$ to mark regions where unphysical oversaturation is likely to occur; this marking directs ConCorrNet’s attention to these sensitive regions, but does not automatically enforce a correction. Instead, ConCorrNet consists of two neural networks that respectively predict the corrective tendencies $\mathrm{d}Q_{\text{fix}}$ and $\mathrm{d}s_{\text{fix}}$. This adaptive methodology allows the model to learn the necessary physical corrections directly from the SPCAM labels for a given atmospheric state, enabling it to reproduce the complex physics of the reference simulation’s condensation processes.
In particular, the correction terms are then applied to the initial tendencies from BasicNet using the humidity mask $\mathrm{Mask}_{\text{h}}$:
$\mathrm{d}Q_{\text{fixed}}$ $=\mathrm{d}Q-\mathrm{Mask}_{\text{h}}\odot\mathrm{d}Q_{\text{fix}}$ (3)
$\mathrm{d}s_{\text{fixed}}$ $=\mathrm{d}s+\mathrm{Mask}_{\text{h}}\odot\mathrm{d}s_{\text{fix}},$ (4)
where $\odot$ denotes element-wise multiplication. Through this mechanism, the physical constraints learned by ConCorrNet are integrated with the predictions from BasicNet. This ensures that the final outputs of the CondensNet model, the tendencies $\mathrm{d}Q_{\text{fixed}}$ and $\mathrm{d}s_{\text{fixed}}$, are physically consistent. Following the methodology of Wang et al. (2022) wang2022stable, the surface precipitation rate is then derived by vertically integrating CondensNet’s final prediction for the water vapor tendency, $\mathrm{d}Q_{\text{fixed}}$. This process provides the necessary moisture source for the land and ocean components of the host GCM, thereby closing the water cycle.
To validate CondensNet and its ConCorrNet’s ability to enforce physical constraints and stabilize simulation, we used six ResMLP models from Wang et al. (2022) wang2022stable recorded as causing unstable simulations. With the weights of these unstable ResMLP models frozen to act as our BasicNet, we trained only the ConCorrNet module. This end-to-end training was guided by a unified loss function on the final, corrected tendencies ($\mathrm{d}Q_{\text{fixed}}$, $\mathrm{d}s_{\text{fixed}}$), ensuring that backpropagated gradients updated only ConCorrNet’s parameters. This experimental design isolates and demonstrates the corrective power of our module (i.e., six corrected cases in Subsection 2.1). Further training specifications are provided in Subsection 4.2.
Ablation studies presented in Supplementary Information section C.2 further validate that correcting both $\mathrm{d}Q$ and $\mathrm{d}s$ is crucial for simulation stability.
4.2 Dataset and training details
@@
The output variables are the corresponding tendencies of water vapor $\mathrm{d}Q$ and dry-static-energy $\mathrm{d}s$ at each vertical level (30 in total), as well as the four radiation fluxes ($SOLS$, SOLLSOLL, SOLSDSOLSD, and $SOLLD$) in which reach to surface.
Notably, in CondensNet, following the traditional column-based parametrization design in GCMs, each neural network instance processes a single atmospheric column independently. During training, column samples from different spatial locations are randomly shuffled, as the network only needs to learn the vertical physical processes within individual columns. When coupled to the host GCM, CondensNet instances operate independently on each column and physics time step. The exchange of mass, momentum, and energy between columns is mediated by the model dynamics, represented through the large-scale tendencies (input variables $\mathrm{d}Q_{\mathrm{l.s.}}$ and $\mathrm{d}T_{\mathrm{l.s.}}$). This design maintains the intrinsic parallel efficiency of parametrizations while retaining the essential horizontal coupling provided by the dynamics.
The inputs consist of vertical profiles of atmospheric state variables ($Q$, $T$), large-scale tendencies ($\mathrm{d}Q_{\mathrm{l.s.}}$, $\mathrm{d}T_{\mathrm{l.s.}}$), and surface conditions ($P_{\mathrm{s}}$, SolinSolin). The outputs include physical tendencies of water vapor ($\mathrm{d}Q$) and dry-static-energy ($\mathrm{d}s$) at each vertical level, along with surface radiation fluxes (summarized in Table 2).
@@
Large-scale water vapor tendency $\mathrm{d}Q_{\mathrm{l.s.}}$ [g/kg/s] 30 144 96
@@
Water vapor tendency $\mathrm{d}Q$ [g/kg/s] 30 144 96
@@
The basic neural networks, BasicNet, is a pre-trained Residual Multilayer Perceptron (ResMLP) that predicts basic tendencies of water vapor ($\mathrm{d}Q$) and dry-static-energy ($\mathrm{d}s$). It contains 7 residual blocks (14 hidden layers in total) and 2 separate ResMLP neural networks – one for predicting $\mathrm{d}Q$ and one for $\mathrm{d}s$. Each ResMLP module has 512 neurons for each hidden layer and uses ReLU activation functions.
@@
The condensation correction network (ConCorrNet) is also a ResMLP designed to adjust the predictions of BasicNet to enforce physical constraints related to water vapor saturation. ConCorrNet architecture includes 6 residual blocks, each containing 2 fully connected layers with 512 neurons, resulting in a total depth of 12 hidden layers. We selected the sigmoid activation function based on its superior convergence performance observed in preliminary experiments.
In terms of training, we focus here on the definition and interpretation of the loss functions used to train CondensNet. BasicNet and ConCorrNet can be optimized either jointly or in a two-stage scheme in which BasicNet is pretrained and then frozen while training ConCorrNet. In this work we adopt the latter for faster convergence; the mathematical objectives below apply to both training protocols.
We minimize a single supervised loss on the final, mask-corrected tendencies – $\mathrm{d}Q_{\text{fixed}}$ and $\mathrm{d}s_{\text{fixed}}$ – as defined in Equation 3 and  4, namely:
$L_{\text{CondensNet}}=\frac{1}{N}\sum_{i=1}^{N}\Big[(\mathrm{d}Q_{\text{fixed},i}-\mathrm{d}Q_{\text{label},i})^{2}+(\mathrm{d}s_{\text{fixed},i}-\mathrm{d}s_{\text{label},i})^{2}\Big].$ (5)
Here, $N$ denotes the number of training samples (grid columns), and $\mathrm{d}Q_{\text{fixed}}$ and $\mathrm{d}s_{\text{fixed}}$ are the masked, physically corrected outputs of CondensNet. This formulation of the minimization problem allows the backpropagation of gradients to both BasicNet and ConCorrNet; the humidity mask then determines when and where ConCorrNet is active, and it needs to learn correcting BasicNet unphysical predictions.
To better understand the training strategy adopted, it is useful to express the minimization problem as the sum of two interpretable terms. First, the standard supervised loss for BasicNet
$L_{\text{BasicNet}}=\frac{1}{N}\sum_{i=1}^{N}\Big[(\mathrm{d}Q_{i}-\mathrm{d}Q_{\text{label},i})^{2}\quad+(\mathrm{d}s_{i}-\mathrm{d}s_{\text{label},i})^{2}\Big].$ (6)
Second, the residual-regression objective for ConCorrNet, restricted to the oversaturated region identified by the binary mask $\mathrm{Mask}_{\text{h}}$ (with cardinality $N_{m}$):
@@
$L_{\text{ConCorrNet}}$ $=\frac{1}{N_{m}}\sum_{i\in\mathrm{Mask}_{\text{h}}}\Big[\left(\mathrm{d}Q_{\text{fix},i}-(\mathrm{d}Q_{i}-\mathrm{d}Q_{\text{label},i})\right)^{2}$
$\quad+\left(\mathrm{d}s_{\text{fix},i}-(\mathrm{d}s_{\text{label},i}-\mathrm{d}s_{i})\right)^{2}\Big].$ (7)
Intuitively, if we interpret the humidity mask as a binary gate: when $\mathrm{Mask}_{\text{h}}=1$ ConCorrNet learns to correct the BasicNet predictions, while when $\mathrm{Mask}_{\text{h}}=0$ ConCorrNet is inactive. Accordingly, the implemented loss in Equation 5 back-propagates gradients to both heads wherever $\mathrm{Mask}_{\text{h}}=1$, and to BasicNet everywhere. Details are provided in Supplementary Information section C.3.
@@
4.3 Climatological Processing